OpenStudy (anonymous):

A person flies a plane against the wind for ninety miles and gets there in 15 minutes. The plane flies back with the wind this time for ninety miles and gets there in nine minutes. How much is the planes speed and winds speed?

OpenStudy (anonymous):

you will need to come up with 2 equation. I will come up with 1 for you. flying againts winds ---> plane speed - windspeed let v = plane speed w= wind speed 15 minutes= 1/4 hour speed * time = distance (v-w) * 1/4hour =